# The collision and merger products of stars do not look alike: A magnetohydrodynamics comparison

### v1

@@ 3.5 Magnetic field strengths
This amplification in $E_{\mathrm{B}}$ provides little information regarding the longevity of magnetic fields. If field lines have many small-scale reversals, magnetic reconnection can diminish field strengths. A large-scale ordered magnetic field, in addition to differential rotation, is necessary in generating a large-scale dynamo that is required for its long-term stability (e.g., R. Pakmor et al., 2024b). The center-left column of Figure 6 shows the edge-on slices of the azimuthal (toroidal) component of the magnetic field, $B_{\phi}$, for the 8 + 7 collision and merger products at the end of the simulations. Whereas the $b=0.25$ collision product shows many small-scale reversals in $B_{\phi}$, the $b=0.50$ collision product and the merger product do not. Small-scale reversals are also observed in all $b=0.25$ collision products, and large-scale order is seen in $b=0.50$ collision products (especially in the outer regions) and merger products of different masses. Crucially, this implies that grazing (large impact parameter) collisions, akin to mergers, result in stars that can potentially generate large-scale dynamos and sustain magnetic fields.
3.6 Magnetized bipolar outflows
A quantitative indicator of the significance of magnetic fields in gas dynamics is the plasma beta parameter, defined as $\beta_{\mathrm{mag}}=P/P_{\mathrm{mag}}$. Here, $P$ is the thermal pressure (gas + radiation) and $P_{\mathrm{mag}}=|\mathbf{B}|^{2}/8\pi$ is the magnetic pressure. Thus, $\beta_{\mathrm{mag}}\ll 1$ suggests that magnetic pressure dominates gas dynamics and vice versa. The center-right column of Figure 6 shows the magnetic $\beta_{\mathrm{mag}}$ of the surrounding material for the 8 + 7 collision and merger products at the end of the simulations. $\beta_{\mathrm{mag}}>10^{4}$ ($\gg 1$) for the $b=0.25$ collision product, indicating that its magnetic field plays no role in gas dynamics. $\beta_{\mathrm{mag}}\sim 10$$100$ for the $b=0.50$ collision product, notably higher and implying that its magnetic field plays a minor role in gas dynamics. However, the non-axial and axial regions, respectively, of the merger products have $\beta_{\mathrm{mag}}\sim 1$$100$ and $\sim 0.01$11, underscoring the importance of magnetic fields in governing gas dynamics, most notably in driving bipolar axial outflows.
The rightmost column of Figure 6 shows the edge-on slices of radial velocities $v_{\mathrm{rad}}$ of the surrounding material for the 8 + 7, $b=0.25$ collision and merger products at the end of the simulations. We see a stark disparity in their radial velocities: the collision products do not show any directed outflows ($v_{\mathrm{rad}}<100\,\mathrm{km}\,\mathrm{s}^{-1}$), while the merger product shows distinct bipolar axial outflows with $v_{\mathrm{rad}}>300\,\mathrm{km}\,\mathrm{s}^{-1}$. Bipolar outflows are present in merger products of stars of different masses and absent in all collision products. These outflows have lower densities than the surrounding material (see Figure 3), and have $v_{\mathrm{rad}}\gtrsim 400\,\mathrm{km}\,\mathrm{s}^{-1}$ in most cases. The 8 + 7 and 10 + 5 (not shown) mergers have the most and least prominent outflows, respectively.
Many MHD simulations of stellar interactions (e.g., K. Kiuchi et al., 2024; J. Morán-Fraile et al., 2024; R. Pakmor et al., 2024b; P. A. Ondratschek et al., 2022) have shown magnetically-driven bipolar outflows of plasma. Although the exact origin of outflows is unclear, they require a strong magnetic field at the center and a long-lived interaction to develop and collimate gas along the rotation axis. The difference in the bipolar outflows between collision and merger products may partly be explained by the timescale of interaction. A collision occurs in a few impulsive stellar passages, ranging from $\sim 10\,\mathrm{hr}$ when $b=0.25$ to $\sim 80$$100\,\mathrm{hr}$ when $b=0.50$, which does not seem to be long enough for magnetized outflows to develop and collimate. In contrast, stars undergoing a binary merger interact for much longer timescales (see Figure 7), ranging from $\sim 200\,\mathrm{hr}$ for the 10 + 5 merger to $\sim 600\,\mathrm{hr}$ for the 8 + 7 merger [^6] Given a constant inspiral velocity, stars with a more skewed mass ratio merge more quickly than those with a near-equal mass ratio., during the accelerated inspiral phase. In reality, their timescales of interaction would be much longer than the orbital timescales, which would provide sufficient time for outflows to develop and strengthen.
4 Summary and conclusions
@@
Shapes of density contours Ellipsoidal More flattened ellipsoidal Peanut-shaped
@@
The aim of this work is to compare the differences and similarities between two types of stellar interactions: collisions and mergers of two MS stars. Employing the 3D moving-mesh code AREPO, we conducted 14 high-resolution MHD simulations of non-rotating MS stars realistically modeled using the 1D stellar evolution code MESA: 8 collisions and 6 mergers for stellar masses ranging from $5\,\mathrm{M}_{\odot}$ to $10\,\mathrm{M}_{\odot}$. In the case of collisions, we set the impact parameters $b=r_{\mathrm{p}}/r_{\mathrm{tot}}$ of the initial parabolic orbits at $b=0.25$ (closer to head-on) for 6 simulations and at $b=0.50$ (farther from head-on) for 2 simulations. In the case of mergers, we assumed initially circular binary orbits that finally merge due to artificially-accelerated inspiral. Our main results are detailed below, while Table 3 provides a concise summary of the major differences between collision and merger products.
[…]
@@
  • Collision and merger products are fast and differentially rotating due to the conversion of orbital angular momenta into spin angular momenta. $b=0.50$ collision products and merger products rotate up to 0.80.80.90.9 of break-up velocity, while $b=0.25$ collision products rotate slower at 0.50.50.60.6 of break-up velocity.
  • As a result of rotation, collision and merger products are flattened. The degree of flattening is higher for faster-rotating remnants. Unlike collision products, the edge-on density contours of merger products are peanut-shaped, indicating the presence of low-density gas along the rotation axes.
@@
  • During the course of stellar interactions, turbulent mixing results in exponential amplifications of magnetic field energies $E_{\mathrm{B}}$ by 991212 orders of magnitude. $b=0.25$ collision products have systematically lower $E_{\mathrm{B}}$, by 1133 orders of magnitude, than the $b=0.50$ collision products and merger products.
@@
  • Magnetized, bipolar axial outflows are launched only from merger products, driven by strong magnetic fields, whereas we find no sign of outflows in collisions. The plasma beta parameters of these outflows are $\beta_{\mathrm{mag}}\sim 10^{-1}$$10^{-2}$ and outflow velocities can exceed $v_{\mathrm{rad}}\sim 300$$400\,\mathrm{km}\,\mathrm{s}^{-1}$.
@@
PV thanks Robert Mathieu and Selma de Mink for helpful discussions. PV thanks Bart Ripperda and Norman Murray for providing access to their computational time on Compute Canada’s Niagara cluster. Most of the simulations presented in this paper were carried out using computational resources (and/or scientific computing services) at the Max-Planck Computing & Data Facility. {contribution}PV performed the simulations in this paper and led the analysis and writing. All authors contributed to the project planning, interpretation of the analysis, and writing.

### v2

@@ III.5 Magnetic fields
Lastly, we explore the longevity of magnetic fields by examining the field morphology. A magnetic field with many small-scale reversals can diminish in strength over time owing to magnetic reconnection. In contrast, a field with large-scale order, in addition to differential rotation, may be powered by a large-scale dynamo and can indicate longevity (e.g., Kiuchi et al., 2024; Pakmor et al., 2024b; Schneider, 2025). The second column of Figure 7 shows the edge-on slices of the toroidal component of the magnetic field, $B_{\phi}$, for the 8 + 7 collision and merger products at the end of the simulations. Whereas the $b=0.25$ collision product shows many small-scale reversals in $B_{\phi}$, the $b=0.50$ collision product and the merger product do not. Small-scale reversals are also observed in all $b=0.25$ collision products, and large-scale order is seen in $b=0.50$ collision products (especially in the outer regions) and merger products of different masses. Crucially, this implies that grazing (large impact parameter) collisions, akin to mergers, result in stars that can potentially generate large-scale dynamos and sustain magnetic fields.
Figure 10: Profiles of average radial velocities in the bipolar axial regions (see text for definition) $v_{\mathrm{rad,ax}}$ as functions of vertical (axial) distances $|z|$. The left and right panels represent collision and merger products, respectively, of different masses. Merger products have clear axial outflows with velocities $v_{\mathrm{rad,ax}}\sim 100$$300\,\mathrm{km}\,\mathrm{s}^{-1}$ while collision products do not.
@@
Shapes of density contours Oblate More flattened oblate Peanut shaped
@@
III.6 Magnetized bipolar outflows
A quantitative indicator of the significance of magnetic fields in gas dynamics is the plasma beta parameter, defined as $\beta_{\mathrm{mag}}=P/P_{\mathrm{mag}}$. Here, $P$ is the thermal pressure (gas + radiation) and $P_{\mathrm{mag}}=|\bm{B}|^{2}/8\pi$ is the magnetic pressure. Thus, $\beta_{\mathrm{mag}}\ll 1$ suggests that magnetic pressure dominates gas dynamics and vice versa. The third column of Figure 7 shows the magnetic $\beta_{\mathrm{mag}}$ of the surrounding material for the 8 + 7 collision and merger products at the end of the simulations. $\beta_{\mathrm{mag}}>10^{4}$ ($\gg 1$) for the $b=0.25$ collision product, indicating that its magnetic field plays no role in gas dynamics. $\beta_{\mathrm{mag}}\sim 10$$100$ for the $b=0.50$ collision product, notably higher and implying that its magnetic field plays a minor role in gas dynamics. However, the nonaxial and axial regions of the merger products have $\beta_{\mathrm{mag}}\sim 1$$100$ and $\sim 0.01$11, respectively, underscoring the importance of magnetic fields in governing gas dynamics, most notably in driving bipolar axial outflows.
Figure 10 shows the average radial velocities in the bipolar axial regions for all collision and merger products at the end of the simulations. The axial regions are defined as bicones of $30^{\circ}$ centered on the CoMs of the remnants and oriented perpendicular to the orbital planes. We see a stark disparity in their radial velocities: the collision products do not show strong, directed outflows ($v_{\mathrm{rad,ax}}<50\,\mathrm{km}\,\mathrm{s}^{-1}$), while the merger product shows distinct bipolar axial outflows with $v_{\mathrm{rad,ax}}\sim 100$$300\,\mathrm{km}\,\mathrm{s}^{-1}$. This is also evident from the fourth column of Figure 7, which shows the edge-on slices of radial velocities $v_{\mathrm{rad}}$ (in all directions) of the surrounding material for the 8 + 7 collision and merger products. In addition, the outflow velocities are higher farther away from the merger products, large enough to escape them and become unbound; the magnetized outflows carry away $\sim 5\%$$10\%$ of the total magnetic energies of the merger products.
Many MHD simulations of stellar interactions (e.g., Ondratschek et al., 2022; Kiuchi et al., 2024; Morán-Fraile et al., 2024; Pakmor et al., 2024b; Vetter et al., 2025) have shown magnetically driven bipolar outflows of plasma. Although the exact origin of outflows is unclear, they require a strong magnetic field at the center and a long-lived interaction to develop and collimate gas along the rotation axis. We speculate that the difference in the bipolar outflows between collision and merger products may partly be explained by the timescale of interaction. A collision occurs in a few impulsive stellar passages, ranging from $\sim 10\,\mathrm{hr}$ when $b=0.25$ to $\sim 80$$100\,\mathrm{hr}$ when $b=0.50$, which does not seem to be long enough for magnetized outflows to develop and collimate. In contrast, stars undergoing a binary merger interact for much longer timescales (see Figure 8), ranging from $\sim 200\,\mathrm{hr}$ for the 10 + 5 merger to $\sim 600\,\mathrm{hr}$ for the 8 + 7 merger [^8] Given a constant inspiral velocity, stars with a more skewed mass ratio merge more quickly than those with a near-equal mass ratio., during the accelerated inspiral phase. In reality, their timescales of interaction would be much longer than the orbital timescales, which would provide sufficient time for outflows to develop and strengthen. However, a more detailed investigation of the physical origin of outflows is beyond the purview of this study.
IV Summary and conclusions
@@
The aim of this work is to compare the differences and similarities between two types of stellar interactions: collisions and mergers of two MS stars. Employing the 3D moving-mesh code AREPO, we conducted 14 high-resolution MHD simulations of nonrotating MS stars realistically modeled using the 1D stellar evolution code MESA: 8 collisions and 6 mergers for stellar masses ranging from $5\,\mathrm{M}_{\odot}$ to $10\,\mathrm{M}_{\odot}$. In the case of collisions, we set the impact parameters $b=r_{\mathrm{p}}/r_{\mathrm{tot}}$ of the initial parabolic orbits at $b=0.25$ (closer to head-on) for 6 simulations and at $b=0.50$ (farther from head-on) for 2 simulations. In the case of mergers, we assumed initially circular binary orbits that finally merge owing to artificially accelerated inspiral. Our main results are detailed below, while Table 3 provides a concise summary of the major differences between collision and merger products.
  • A direct consequence of stellar collisions and mergers is the chemical mixing of stars, resulting in hydrogen in the initial stellar envelopes being mixed into the new stellar cores and helium in the initial stellar cores being mixed into the new stellar envelopes. Although the cores are rejuvenated in both types of interactions, the exact core hydrogen fractions in the merger products can be $3\%$$10\%$ higher than the corresponding collision products in some cases. Thus, their MS lifetimes are also longer by a similar fraction, which may affect future stellar evolution.
@@
  • Collision and merger products are fast and differentially rotating owing to the conversion of orbital angular momenta into spin angular momenta. The $b=0.50$ collision products and merger products rotate up to 0.80.80.90.9 of breakup velocity, while $b=0.25$ collision products rotate slower at 0.50.50.60.6 of breakup velocity.
  • As a result of rotation, collision and merger products are flattened. The degree of rotational flattening in the collision products is higher for larger impact parameter ($b=0.50$) encounters. In comparison, the density contours of merger products are peanut shaped, with a dearth of material in the polar regions.
@@
  • During the course of stellar interactions, turbulent mixing results in exponential amplifications of magnetic field energies $E_{\mathrm{B}}$ by 991212 orders of magnitude assuming a small initial field. The $b=0.25$ collision products have systematically lower $E_{\mathrm{B}}$, by 1133 orders of magnitude, than the $b=0.50$ collision products and merger products.
@@
  • Magnetized, bipolar axial outflows are launched only from merger products, driven by strong magnetic fields, whereas we find no sign of outflows in collisions. The plasma beta parameters of these outflows are $\beta_{\mathrm{mag}}\sim 10^{-1}$$10^{-2}$, with outflow velocities of up to $v_{\mathrm{rad}}\sim 300\,\mathrm{km}\,\mathrm{s}^{-1}$.
@@
P.V. thanks Robert Mathieu and Selma de Mink for helpful discussions. P.V. thanks Bart Ripperda and Norman Murray for providing access to their computational time on Compute Canada’s Niagara cluster. Most of the simulations presented in this paper were carried out using computational resources (and/or scientific computing services) at the Max-Planck Computing & Data Facility.
{contribution}
PV performed the simulations in this paper and led the analysis and writing. All authors contributed to the project planning, interpretation of the analysis, and writing.